# Sketch-Plan-Generalize: Continual Few-Shot Learning of Inductively Generalizable Spatial Concepts

### v2

@@ 4 Technical Approach: Learning Inductive Spatial Concepts
The progress exploration of the program space can be summarised as follows:
$H_{S}^{*}\leftarrow Sketch(\Lambda\,;\,{\theta_{S}})\,;\,H_{P}^{*}\leftarrow Plan(S_{1}..S_{g},H_{S}^{*};\,{\theta_{P}})\,;\,H_{G}^{*}\leftarrow Generalize(H_{P}^{*},\,H_{S}^{*};\,{\theta_{G}})$ (3)
The concept library $\mathcal{L}$ is initialized with primitive visual and action concepts. Upon acquiring a new inductive concept $H_{G}^{*}=H^{*}$, we update our library accordingly: $\mathcal{L}\leftarrow\mathcal{L}\cup{H^{*}}$. Next, we detail each of the three steps mentioned above. Fig. 2 illustrates an example of progressive program estimation.
@@
An LLM driven by in-context learning is used to get a program signature (a sketch) for a concept from the natural language instruction. The task sketch is a tree of nested function calls that outlines the function header (name and the parameters) of the inductive concept/program to be learned. A detailed exposition on prompting appears in the Appendix  C.1. The task sketch thus obtained is then grounded on the input scene using a quasi-symbolic visual grounding module akin to [nscl, nsrm-Kalithasan2022LearningNP, wang2023progport]. This module has three key components: (1) a visual extractor (ResNet-34 based) that extracts the features of all objects in the scene, (2) a concept embedding module that learns disentangled representations for visual concepts like green and dice, and (3) a quasi-symbolic executor equipped with pre-defined behaviours such as “filter” to select/ground the objects of interest. For example, grounding the task sketch “Tower (height =3, objects = filter(green, dice))” results in an instantiated function call “Tower(height = 3, objects = [1, 2, 3])” where [1, 2, 3] are the green coloured dice.
@@
Identifying the position $pos_{\theta}(.)$ of the object placement/structure construction is achieved through a notion of head. Essentially, the head represents a cuboidal enclosure in 3D space. The movement of any object involves two key stages: (1) relocating the abstract head to the desired position and (2) picking the target object and placing it at the current position of the head. Conceptually, the movement of the head can be likened to the robot’s cognitive exploration of potential object placements to attain the desired spatial configuration. We define a set of four primitive functions, $\mathcal{A}_{p}$, that manipulate the head. (1) store_head(): remembers the current position of the head in a stack. (2) reset_head(): sets the head to the last remembered position (i.e., to the top of the stack) and pops this position from the stack (3) keep_at_head(objects): places the object corresponding to the index with highest value in objects (a list of values) at the current location of head and (4) move_head(direction): A neural operator that takes the head and predicts the new location of the head with respect to the given direction. The move_head function is trained using a corpus of pick and place instructions, such as “move the green object to the right of the red cube” similar to the approach in [nsrm-Kalithasan2022LearningNP]. When an object is passed as the parameter, the move_head operator is overloaded to assign the head to the location of the object (also referred to as assign_head(object)).
@@
The state is defined by bounding boxes (including the depth of the center) and visual attributes of all the objects that are present on the table. We assume that the head is initially assigned to the position where the structure needs to be constructed. For each learned inductive concept <cpt>, we define a macro-action “Make_<cpt>(size)" that executes the corresponding program with the given size argument, resulting in the construction of the desired inductive concept. Thus, the action space $\mathcal{A}$ is the union of primitive actions $\mathcal{A}_{p}$ and compound/macro-actions $\mathcal{A}_{c}$. All actions except keep_at_head(objects) and macro actions yield a reward of zero, while for the latter, the Intersection over Union (IoU) between the attained state and the expected state in the demonstration is provided as a reward. An MCTS procedure is performed to find a plan that maximizes the reward. The search outputs a sequence of grounded actions for an instantiation of the given inductive concept by the task arguments.
@@
MCTS that searches for a plan only in terms of primitive actions may not be generalizable due to lack of modularity  C.3. The use of macro-actions in the search ensures that the plan $H_{p}^{*}$ for a given demonstration is concise, modular, and easily generalizable. This can be seen as a form of regularization in terms of the length of concept description by making the prior $P(H)\propto|H|^{-\alpha}$ (where $\alpha$ > 0) in equation (2) $H^{*}=\arg\min_{H\in\mathcal{H}}\left[\mathtt{Loss}(\{S_{1}..S_{n}\},H(\Lambda))+\alpha\log|H|\right]$. However, as the action space expands with the learning of more concepts, the search becomes slower, necessitating the pruning of the search space. To avoid searching over the size parameter in macro-actions, we greedily select the smallest size that achieves the maximum average reward from the current state. Additionally, to prune primitive actions, we train a reactive policy $\pi_{neural}$ which, given the current state $\tilde{s_{t}}$ and the next expected state ${s_{t+1}}$ (from the demonstration), outputs one of the primitive actions $a_{t}^{*}\in\mathcal{A}_{p}$. Consequently, the effective branching factor of the search is reduced to $|\mathcal{A}_{c}|+1$. Thus, our MCTS algorithm is modular through the hierarchical composition of learned concepts and efficient through action space pruning, and is referred to as MCTS+L+P. Further details regarding modifications in MCTS are given in the Appendix  A.4.
@@
Leveraging the code generation and pattern matching abilities of LLMs [pattern] we use GPT-4 to distil out a general Python program from the sequence of grounded actions as determined by MCTS+L+P. The learnt program is incorporated in the concept library, $\mathcal{L}$, for modular reuse in subsequent learning tasks. Additional details, prompting mechanism and use of learnt programs in the search step of future learning tasks are described in Appendix  C.2,  A.4. We take a curriculum learning approach beginning from learning of primitive actions and visual attributes, followed by structures of increasing complexity. Additional details about learning of inductive concepts via curriculum, architecture details, and learning from multiple demonstrations are in Appendix  C.2, Fig.  7,  A.5.
@@ 5 Evaluation Setup
Data sets. Evaluation data set was created with a simulated Franka Emika Robot Manipulator constructing arrangement of blocks of uniform size with varied type/appearance (e.g., cubes, dice, lego etc.). The scene is observed with a visual and depth sensor with object masks extracted from a standard segmentation procedure. A total of $15$ structure types are considered including towers, staircases, X (cross), arch-bridge. The state and action sequences for attaining these structures forms the demonstration data fro training, details in Appendix  10.
Figure 3: Two illustrative examples of complex spatial structures from our dataset with a hierarchical composition of inductively defined structures.
The training data for each concept consists of $3$ demonstration per structure, of size (num. of inductive steps) $3-5$ with up to $20$ objects. We form three evaluation data sets, each consisting of simple structures (different size instances of a concept) and complex structures composed of simpler concepts (e.g., staircase consists of towers as substructure). Dataset I includes contains structures ranging in sizes $3-5$. Dataset II reverses the linguistic labels (e.g., the “tower" becomes “rewot"). Dataset III includes concepts of larger sizes ($4-7$).
Alternative approaches. We compare our model with four baselines spanning two approaches:
@@
(1) Purely-Neural: An end-to-end neural model inspired by StructDiffusion, [struct-diffusion] that treats structure construction as a rearrangement problem. We consider two variations of the model 1.1) Struct-Diff (SD): End-to-end approach without any additional supervision regarding which objects need to be moved. 1.2) With-Grounder (SD+G): Similar to 1.1) except that we assume a perfect object selector/grounder that identifies the relevant set of objects which are to be moved.
@@
(2) Pre-trained models that directly output symbolic programs: 2.1) LLM for Scene-Graph Reasoning: pre-trained language models (GPT-4) that generate Python programs from instructions which would describe the given demonstration. To help LLM understand the underlying structure, symbolic spatial relationships (e.g., left(a,b)) between objects in the demonstration are also provided to the LLM. For this baseline we further assume absence of distractor objects in the scene. 2.2) Vision Language Model (GPT-4-V): Similar to 2.1 but has the ability to take input demonstration as images. For learning program of a new inductive concept we give the demonstration to the VLM in the form of $\Lambda,(S_{1}..S_{g})$. Additional details on prompting method in Appendix C.6111Experiments were also conducted with open source code-generation LLM such as CodeLlama (70Bq), Due to significantly poorer performance w.r.t. GPT-4, it was removed as the primary LLM baseline..
@@
Model variants. We implement three variants of the MCTS search to perform a grounded plan search over the action space $\mathcal{A}$: (i) MCTS+P+L: Our approach as described in section  4.2, that uses the learnt concepts from $\mathcal{L}$ as macro actions (e.g., Make_Tower(3, objects) $\in\mathcal{A}_{c}$) (L). Further it performs pruning of $\mathcal{A}_{p}$ using $\pi_{\text{neural}}$ (P), (ii) MCTS-P+L: Our approach without neural pruning and (iii) MCTS+P-L: No access to library of concepts during continual learning and thereby lacks ability to use macro actions. This method greedily selects the action from $\mathcal{A}_{p}$ as given by $\pi_{\text{neural}}$. We provide more details about the 3 methods in Appendix  D.4.
@@
Metrics. We adopt the following metrics to evaluate our models: (i) Program Accuracy: A binary score obtained through human evaluation. (ii) Target Construction IoU: Intersection over Union (2D-IoU) between the bounding boxes (excluding depth). (iii) Target Construction Loss: Mean Squared Error (MSE) loss between bounding boxes (including depth). Note that both IoU and MSE are computed invariant to order between objects with same set of attributes.
@@ 6 Results
Our experiments evaluate the following (i) How does our model perform as compared to baselines in terms of concept learning and execution ability (In-Distribution) ? (ii) How does our model perform in comparison to baselines in ability to generalize to concept instances larger than those seen during training (Out-of-Distribution) ? (iii) How robust and efficient is our learning pipeline ? (iv) How can the acquired concepts be used in complex reasoning and planning tasks ?
6.1 Effectiveness of Concept Learning
Tables  2 and  2 compares the the program accuracy [^2] Purely neural approach is marked as NA due to absence of program representations. and IoU/MSE metrics on attaining the final state of the proposed and the baseline approaches in the in-distribution setting. The proposed SPG approach significantly improves over other approaches. We make the following observations. First, for complex compositional structures, the accuracy of the pre-trained models is poor (zero), indicating their inability to reason over numerous and complex spatial relations inherent in structures (e.g., staircases, bridge, x-shaped etc.). Additionally, we note that program inference via an LLM (GPT4) shows improved performance learning simple structures (e.g., rows or columns) using the tokenized language description of the scene compared to a VLM (GPT-4V) provided with a visual input (reflecting in better IoU/MSE scores). However, we observe that for complex structures, the performance of VLM (GPT-4v) is better than an LLM (GPT) indicating the difficulty in complex spatial reasoning inherent in interpreting complex concepts. Purely neural approach that is completely grounded in the demonstration performs much better on complex structures as compared to pre-trained foundation models. However our approach still outperforms it and is more data efficient.
Table 1: Program Accuracy (PA)
@@ 6.1 Effectiveness of Concept Learning
6.2 Generalization Performance
Table  4 compares the generalization performance on Dataset III [^3] Full table with standard-errors reported in Appendix,  8. The proposed model outperforms other approaches. Further, we observe a relative decrease in 2D-IoU being 7.27% for simple and 5.74% for complex structures. In contrast Struct-Diffusion with grounder shows a relative decrease of 63.25% on simple structures and 74.72% on complex structures; highlighting their inability to generalize inductively. Pre-trained models on the other hand have much higher relative decrease in performance for complex structures (53.87% and 41.64%), which is attributed to their inability generate the correct program that can generalize inductively to unseen data. Table 9 in Appendix quantifies relative performance decrease values as compared to Dataset I.
Table 3: Out-of-Distribution Performance
@@ 6.2 Generalization Performance
IoU MSE (1e-3) IoU MSE (1e-3)
SPG(Ours) 0.89 0.43 0.80 1.49
GPT-4V 0.58 13.2 0.29 10.9
GPT-4 0.78 5.51 0.13 19.1
SD+G 0.27 6.21 0.15 14.2
SD 0.24 6.86 0.15 11.6
Table 4: Perf. on Dataset II: Names Reversed
@@
IoU MSE (1e-3) IoU MSE (1e-3)
SPG(Ours) 0.86 1.74 0.78 3.93
GPT-4 0.78 3.16 0.00 22.73
GPT-4V 0.71 3.92 0.09 21.29
6.3 Robustness and Efficiency Analysis
6.3.1 Reliance on pre-trained Knowledge vs. Demonstration.
Next, we evaluate the degree to which concept learning relies on background knowledge vs. observed action sequences for its construction. We compare pre-trained models against our approach by learning programs on another training dataset differing only in terms of the names of the structures (names being reversed). Table  10 in the Appendix gives the corresponding program accuracies. Table  4 indicates the corresponding IoU and MSE, with our approach again showing the best performance [^4] For the IoU/MSE values along with standard errors refer to table  11 in Appendix.. The relative decrease in performance for our approach (12%) is lower than GPT-4(14%) and much lower than GPT-4-V(30%) (in terms of program accuracy). The poorer generalization of pre-trained models can be attributed to over-reliance prior knowledge and failing to effectively use demonstrations.
6.3.2 Analysis of MCTS Variants for Concept Learning

### v3

@@ 5 Learning Inductive Concepts from Demonstrations
Formally, the factored exploration of the program space for a demonstration is performed as:
$H_{S}^{*}\leftarrow Sketch(\Lambda\,;\,{\theta_{S}})\,;\,H_{P}^{*}\leftarrow Plan(S_{1}..S_{g},H_{S}^{*};\,{\theta_{P}})\,;\\ \,H_{G}^{*}\leftarrow Generalize(H_{P}^{*},\,H_{S}^{*};\,{\theta_{G}})$ (3)
Here, $\theta_{S}$, $\theta_{P}$ and $\theta_{G}$ are the learnable parameters (including hyperparameters) of the Sketch, Plan and Generalize functions, respectively. The concept library $\mathcal{L}$ is initialized with primitive visual and action concepts. Upon acquiring a new inductive concept $H_{G}^{*}=H^{*}$, we update our library accordingly: $\mathcal{L}\leftarrow\mathcal{L}\cup{H^{*}}$. An example is provided in Appendix Sec. A.5. Fig. 2 illustrates an example of progressive prog. estimation. Next, we detail each of these three steps.
@@
An LLM driven by in-context learning is used to get a program signature (a sketch) for a concept from the natural language instruction. The task sketch is a tree of nested function calls that outlines the function header (name and the parameters) of the inductive concept/program to be learned. Detailed exposition on prompting appears in the app.  C.1.
The task sketch thus obtained is then grounded on the input scene using a visual grounding module akin to (Mao et al., 2019; Kalithasan et al., 2023; Wang et al., 2023c). This module has three key components: (1) a visual extractor (ResNet-34 based) that extracts the features of all objects in the scene, (2) a concept embedding module that learns disentangled representations for visual concepts like green and dice, and (3) an executor equipped with pre-defined behaviours such as “filter” to select/ground the objects of interest. For example, grounding the task sketch “Tower (height =3, objects = filter(green, dice))” results in an instantiated function call “Tower(height = 3, ObjectSet)” where ObjectList is an ordered list of the green coloured dice. Since it is now grounded in the initial scene of the demonstration, the task sketch corresponds to a particular instance of the concept demonstrated in the given demonstration.
@@
Primitive Actions. Constructing complex structures involves two steps: (1) identifying or imagining the placement location of an object/structure and (2) picking and placing the object at the imagined location. The position $pos_{\theta}(.)$ for placement is determined using a head, which represents a cuboidal enclosure in 3D space. Conceptually, moving the head is akin to the robot’s cognitive exploration of potential placements to achieve the desired spatial configuration. We define a set of primitive functions, $\mathcal{A}_{p}$, to guide the movement and placement of objects in two ways: (1) move_head(direction): This primitive is a neural operator that moves the abstract head to a desired relative position. It is trained on a corpus of pick-and-place instructions, similar to the approach in (Kalithasan et al., 2023); and (2) keep_at_head(objects): This primitive places the target object from the list objects at the current location of the head.
@@
MCTS Search. We use an object-centric state representation defined by bounding boxes (including the depth of the center) and visual attributes of all the objects that are present on the table. Each previously learned inductive concept <cpt> has an associated macro-action Make_<cpt>(size), that executes the corresponding program for the given size argument, resulting in the construction of the desired concept. Thus, the set of actions $\mathcal{A}$ is the union of primitive actions $\mathcal{A}_{p}$ and compound/macro-actions $\mathcal{A}_{c}$. Intersection over Union (IoU) between the attained state and the expected state in the demonstration is provided as a reward for all macro actions and keep_at_head(objects); all other actions yield zero reward. An MCTS procedure similar to (Khandelwal et al., 2016) (detailed in Appendix A.3) is performed to find a plan (sequence of grounded actions) that maximizes the reward.
Modularity and Scalability. MCTS that searches for a plan only in terms of primitive actions may not be generalizable due to lack of modularity( C.3). The use of macro-actions in the search ensures that the plan $H_{p}^{*}$ for a given demonstration is concise, modular, and easily generalizable. This can be seen as a form of regularization in terms of the length of concept description by making the prior $P(H)\propto|H|^{-\alpha}$ (where $\alpha$ > 0) in equation (2)
$H^{*}=\arg\min_{H\in\mathcal{H}}\left[\mathtt{Loss}(\{S_{1}..S_{n}\},H(\Lambda))+\alpha\log|H|\right]$
@@
However, as the action space expands with the learning of more concepts, the search becomes slower, necessitating the pruning of the search space. To avoid searching over the size parameter in macro-actions, we greedily select the smallest size that achieves the maximum average reward from the current state. Additionally, to prune primitive actions, we train a reactive policy $\pi_{neural}$ which, given the current state $\tilde{s_{t}}$ and the next expected state ${s_{t+1}}$ (from the demonstration), outputs one of the primitive actions $a_{t}^{*}\in\mathcal{A}_{p}$. Consequently, the effective branching factor of the search is reduced from $N*|\mathcal{A}_{c}|+|\mathcal{A}_{p}|$ to $|\mathcal{A}_{c}|+1$, where $N$ is the average number of size-parameter variants. Thus, our MCTS algorithm is modular through the hierarchical composition of learned concepts and efficient through action space pruning, and is referred to as MCTS+L+P. (Detailed in app.  A.3)
@@
Leveraging the code generation and pattern matching abilities of LLMs (Mirchandani et al., 2023), we use GPT-4 to distil out a general Python program from the sequence of grounded actions as determined by MCTS+L+P. The learnt program is incorporated in the concept library, $\mathcal{L}$, for modular reuse in subsequent learning tasks. We take a curriculum learning approach (App.  D.2). Further details & discussion can be found in app. items C.2A.4 and  A.3.
6 Adapting Learned Concepts for Novel Instructions and Constraints
Goal Grounding: A complex Natural Language (NL) instruction is passed to the Parser, which extracts two components: the relevant concepts and the associated constraints. If some of the concept(s) are not present in the concept library $\mathcal{L}$, the system queries for a Human Demonstration, which is then processed via the SPG framework, and the abstracted & generalised concept(s) are appended to $\mathcal{L}$, as in Figure: 3.
Figure 3: Pipeline for using Solvers and LLMs to solve constraints in Instructions over concepts in Library $\mathcal{L}$ and learning additional concepts via SPG if necessary.
Following this, an LLM identifies whether an explicit constraint resolution is required. If yes, the Language Parser isolates the constraint components from the NL Instruction. , and generates a program of constraints that can be solved independently. This bifurcation allows the system to resolve increasingly complex constraints without entangling them with the core structural learning. This is an improvement over directly relying on LLMs to simultaneously learn structures and resolve complex constraints, since LLMs often fail to reason about spatial grounding, visibility, occlusion, and physical stability (Appendix: D.5). Our pipeline modularizes the constraint resolution stage by invoking neural solver modules, external constraint solvers, or LLMs, based on downstream tasks, thereby improving the pure-LLM baseline. We aim to improve the robustness of our pipeline by integrating advanced reasoning capabilities.
Plan Synthesis: We qualitatively demonstrate that the concepts we have acquired can help us to perform goal conditioned planning. These goals may either be direct instances of concepts in $\mathcal{L}$, or constraint-satisfying goals obtained from sec. 6 above. We perform a forward search using the abstract state and action representations to the inferred goal akin to  (Liu et al., 2024; Kalithasan et al., 2024) (detailed in Appendix D.5).
@@ 7 Evaluation Setup
Corpus. A corpus is created using a simulated Robot Manipulator assembling spatial structures on a table-top, viewed by a visual-depth sensor. Demonstration data ($3$ demonstrations per structure, up to $20$ objects per scene) includes RGBD observations of the action sequences (picking and placing of blocks) resulting in the construction of the final assembly using varied block instances and types (cubes, dice, lego etc.).
Figure 4: Illustrative examples of spatial structures from the corpus, showing inductive composition over simpler structures. Details and visualizations in Appendix  12
The scope of concepts and associated evaluation tasks are adapted from closely related works. The staircase and enclosure construction tasks are inspired from from (Silver et al., 2019), adapted to 3D from the original 2D grid world setting. Structures such as boundaries involving repetitive use of columns and rows (w/o explicit joint fastening) are inspired by a robotic assembly data set (Collins et al., 2024). Finally, the arch-bridge and x-shaped patterns are inspired from concept learning works as  (Lake et al., 2015). A total of $15$ structures types are incorporated and are additionally modulated in size/spatial arrangement for generalization evaluation. Three evaluation data sets are formed each with simple structures and complex structures composed of simpler concepts (e.g., staircase consists of towers as substructure). Dataset I and II contain demonstrations constructing structures with $\mathtt{size}(.)\in[3,5]$, where $\mathtt{size}$ is defined in  4. Dataset II reverses the linguistic labels used (e.g., the “tower" in I becomes “rewot" in II) to assess model reliance on pre-training knowledge in presence of new labels for concepts. Dataset III includes concepts of larger size than those in training to test generalization.
Baselines. Four baselines are formed from two alternative approaches as follows.
@@
(1) Purely-Neural: An end-to-end neural model inspired by StructDiffusion (Liu et al., 2023) that treats structure construction as a rearrangement problem. We consider two variations: (1.1) Struct-Diff+Grounder(SD+G): End-to-end approach that assumes a perfect object selector/grounder which identifies the relevant set of objects to be moved. (1.2) Struct-Diff(SD): (1.1) without assuming a grounder.
@@
(2) Pre-trained models that directly output symbolic programs: (2.1) LLMs for Scene-Graph Reasoning: This approach uses a Pre-trained Language Model (GPT-4) to generate Python programs from instructions which describe the given demonstration. Input is provided to the LLM via textual symbolic spatial relationships (e.g., left(a,b)) between objects in the demonstration. This baseline further assumes the absence of distractor objects in the scene. (2.2) Vision Language Model (GPT-4V): Similar to (2.1) but has the ability to take input demonstration as images. For learning the program of a new inductive concept, we give the demonstration to the VLM in the form of $\Lambda,(S_{1}..S_{g})$. Additional details on prompting method in Appendix C.6. Experiments were also conducted with open-source LLMs such as CodeLlama (70Bq), Due to significantly poorer performances w.r.t. GPT-4, GPT-4 was retained as the primary LLM baseline.
@@
Model variants. We implement three variants of the MCTS search to perform a grounded plan search over the action space $\mathcal{A}$: (i) MCTS+P+L: Our approach as described in section 5.2, that uses the learnt concepts (L) from $\mathcal{L}$ in subsequent searches, along with pruning (P) $\mathcal{A}_{p}$ using $\pi_{\text{neural}}$, (ii) MCTS-P+L: Our approach without neural pruning and (iii) MCTS+P-L: No access to library of concepts during continual learning. This method greedily selects the action from $\mathcal{A}_{p}$ as given by $\pi_{\text{neural}}$. For details refer App. D.4.
@@
Metrics. We adopt the following metrics to evaluate our models: (i) Program Accuracy: A binary score obtained through human evaluation. $1$ for constructing the structure fully, 00 otherwise. (ii) Target Construction IoU: Intersection over Union (2D-IoU) between bounding boxes. (iii) Target Construction Loss: Mean Squared Error (MSE) loss over the bounding boxes + depth of the center.
@@ 8 Results
Our experiments evaluate the following questions. Q1: How does our model perform when compared to baselines in terms of concept learning and execution ability (In-Distribution)? Q2: How does our model generalize to concept instances not seen (larger) during training (Out-of-Distribution)? Q3: How robust and efficient is our concept learning pipeline? Q4: How can the acquired concepts be used in for embodied instruction following tasks?
Q1: Concept Learning Accuracy
We compare the program accuracy (Table 1) and the IoU/MSE values (Table 2) of the final states attained by SPG and the baselines w.r.t. the gold states in the in-distribution setting and find that SPG significantly outperforms other approaches. Values for Purely neural approaches are marked NA because Neural Outputs are not physically grounded. We make the following observations: (i) For complex compositional structures, the accuracy of the pre-trained models is poor (zero), indicating their inability to reason over the numerous and complex spatial relations present in these structures. (ii) While program inference via the LLM is better than the VLM for learning simple structures, it is worse for complex structures. This indicates the inherent weakness of the textual descriptions of complex spatial relations present in complex structures. (iii) While the data-intensive purely neural approaches perform much better on complex structures when compared to the pre-trained foundation models, they are still weaker than SPG.
Table 1: Program Accuracy
@@
Q2: Generalization Performance
Table 3, compares the generalization performance on Dataset III for models trained on Dataset I (full table in Appendix,  9). We see that SPG outperforms other approaches. We further consider the relative decrease (R.D.) in performance (2D-IoU) on going from the in-distribution to the out-of-distribution (OOD) setting. We make the following observations: (i) SPG suffers a R.D. of 7.27% for simple and 5.74% for complex structures. (ii) In contrast, the SD+G baseline shows a large R.D. of 63.25% on simple structures and 74.72% on complex structures; highlighting the inability of Purely Neural Models to generalize inductively. (iii) Pre-trained models also have a large R.D. in perf. for complex structures (GPT-4 : 53.87% & GPT4V : 41.64%), which is attributed to their inability to generate the correct program that can generalize inductively to unseen data.
Table 3: OOD Performance. R.Dis the relative decrease in IoU from Table 2. MSE is in 1e-3 units
@@
IoU R.DMSE IoU R.DMSE
SPG(Ours) 0.89 7.27 0.43 0.80 5.74 1.49
GPT-4V 0.58 23.33 13.2 0.29 41.64 10.9
GPT-4 0.78 12.61 5.51 0.13 53.87 19.1
SD+G 0.27 63.25 6.21 0.15 74.72 14.2
SD 0.24 51.84 6.86 0.15 67.67 11.6
Q3: Robustness and Efficiency Analysis
Reliance on pre-trained Knowledge vs. Demonstration. Next, we evaluate the degree to which concept learning relies on prior knowledge vs. the action sequences observed in demonstrations. We compare pre-trained models against our approach by learning programs on Dataset II7), which uses arbitrary names for concepts. This forces all models to rely on demonstration data because there is no real-world knowledge associated with the name of the concept, say "rewot" instead of "tower". Table  4 indicates the corresponding performances, with our approach outperforming others. For the IoU/MSE values along with standard errors refer to app. Table  10. The R.D. in performance (program accuracy w.r.t. Table  1) for simple structures for our approach (12%) is lower than GPT-4 (14%) & much lower than GPT-4V (30%). The poorer generalization of pre-trained models can be attributed to their over-reliance on prior knowledge and failure to effectively incorporate the data from demonstrations. In contrast, SPG better captures the semantics of a novel concept, especially ones whose knowledge may not be available for the LLMs/VLMs at training time.
Table 4: Perf. on Dataset II with Reversed Names. Acc. is Prog. Accuracy, MSE in 1e-3 units
@@
Acc. IoU MSE Acc. IoU MSE
SPG(Ours) 0.88 0.86 1.74 0.78 0.78 3.93
GPT-4V 0.23 0.71 3.92 0.00 0.09 21.29
GPT-4 0.67 0.78 3.16 0.00 0.00 22.73
MCTS Variants for Concept Learning. Figure  5 compares the program accuracy for the three methods of plan search. For the MCTS-L+P method, the program accuracy is expected to be independent of expansion steps as it greedily chooses the action for which $\pi_{\text{neural}}$ gives the highest probability. For the MCTS+L based methods the accuracy increases beyond 0.6 with time, which demonstrates that having a composable library of concepts allows us to learn a much richer class of inductive concepts. MCTS+P+L saturates to a program accuracy of 0.933 in just 4000 expansion steps compared to MCTS+L-P taking 512000 expansion steps, which demonstrates a significant increase in learning efficiency via use of the neural pruner. For very low number of expansions steps (<40) accuracy of MCTS+L based methods is lower than MCTS-L as the former expends expansion steps on UCB exploration (instead of greedy actions).